What is the final volume in milliliters when 0.952L of a 42.3\% (m/v) solution is diluted to 20.2\% (m/v)?
1
Expert's answer
2013-04-12T08:01:23-0400
If 0.952 L contain 42.3% (m/v):
x/0.952 = 0.423.
The mass of a dissolved substance is: 0.423 x 0.952 = 0.403 grams.
So if the original amount of a solution is being diluted to 20.2%, the mass of
a dissolved substance remains the same.
0.403/x = 0.202
x = 0.403/0.202 = 1.99
That is why the final amount of a solution 20.2% (m/v) is 1.99 L.